what is 4937 written in roman numerals?
Found 4 solutions by prabhjyot, JimMarshall, CoolRekt, The Troll Kid: Answer by prabhjyot(165) 
 4937
This can be written as
=4000+900+30+7
=MMMM + CM + XXX + VII
=MMMMCMXXXVII
Basics:
A numeral is a symbol used to represent a number. (Our digits 0-9 are often called Arabic numerals.) Each letter used in Roman numerals stands for a different number:
A smaller number in front of a larger number means subtraction, all else means addition. For example, IV means 4, VI means 6.
You would not put more than one smaller number in front of a larger number to subtract. For example, IIV would not mean 3.
You must separate ones, tens, hundreds, and thousands as separate items. That means that 99 is XCIX, 90 + 9, but never should be written as IC. Similarly, 999 cannot be IM and 1999 cannot be MIM.
I :The numeral one.
II: is two, III: is three. You seldom see IIII as 4, since IV can also mean 4, plus its shorter to write.
V: The numeral 5. IV: is 4, VI: is 6, VII: is 7, VIII: is 8.
X: The numeral 10. IX; is 9, XI: is 11, etc.
L: The numeral 50. XL: would be 40.
C: The numeral 100. Think of Century having a hundred years. C is short for the Latin word Centum.
D: The numeral 500.
M: The numeral 1000.
Sometimes you will see a numeral with a line over it. That means to multiply it by 1000. A numeral V with a line over it means 5000.
Answer by JimMarshall(82)  Answer by CoolRekt(53) 
 MMMMCMXXXVII = 4937
MMMM = 4000
CM = 900
XXX = 30
V = 5
II = 2
4000 + 900 + 30 + 5 + 2 = 4937
MMMM + CM + XXX + V + II = MMMMCMXXXVII = 4937
CoolRekt
Answer by The Troll Kid(22) 
| |
|